# Sphere Packing on a Quantum Computer for Chromatography Modeling

### v1

@@ Abstract
Protein chromatography is an important technique in biopharmaceutical manufacturing that separates proteins by filtering them through a tightly packed column of gels. Tighter packings yield better protein separation. To this end, we model chromatography as sphere packing, formulating three models, each with increasing complexity. The first, homogeneous circle packing, is recast as maximum independent set and solved by the Quantum Approximate Optimization Algorithm on a quantum computer. The second, heterogeneous circle packing, is formulated as a graphical optimization problem and solved via classical simulations, accompanied by a road map to a quantum solution. An extension to the third, heterogeneous sphere packing, is formulated mathematically in a manner suitable to a quantum solution, and detailed resource scaling is conducted to estimate the quantum resources required to simulate this most realistic model, providing a pathway to quantum advantage.
@@ I Introduction
As quantum computational power continues to grow rapidly, it is paramount that industries which rely heavily on computation start to plan today for tomorrow’s quantum advantage. One such industry is the biopharmaceutical industry, which uses massive amounts of classical computation to aid everything from drug discovery to operational optimization. In this work, we investigate the utility of quantum computing to aid protein chromatography, in which spherical gels are packed tightly into a cylindrical column through which a protein mixture is filtered. The tighter the gels are packed, the better the proteins are separated. This problem is modeled as bounded sphere packing, in which spheres are packed to minimize empty space. Bounded sphere packing is further cast as a graphical optimization problem amenable to quantum solutions.
We break the problem of sphere packing into three levels of increasing complexity: homogeneous circle packing, heterogeneous circle packing, and heterogeneous sphere packing. We were able to solve an example of homogeneous circle packing on real quantum hardware that exists today. Details of the hardware experiment include discussions on hyperparameter optimization, compilation, and noisy simulation. Evidence of parameter concentration is also included. Numerical results for heterogeneous circle packing via classical computation were executed, and the problem was formulated in a manner suitable for quantum computation. Finally, heterogeneous sphere packing is shown to be an extension of the heterogeneous circle packing formulation, and resource estimation is performed to estimate the quantum resources required to solve it. We will argue that, while classical computers can approximately solve all three formulations today, the difficulty in doing so scales exponentially with increasing problem size and formulation complexity. In contrast, while quantum computers can only solve the simplest formulation today, there is good evidence that in time, they will be able to solve all three levels of complexity, each with polynomially scaling difficulty.
@@
Figure 1: Three complexity levels of sphere packing for modeling chromatography: homogeneous circle packing, heterogeneous circle packing, and heterogeneous sphere packing. Also shown are the difficulty levels of classical and quantum solutions over time for the three complexity levels.
I.1 Protein Chromatography
@@
Figure 2: Schematic of chromatography for protein filtration. Different proteins filter through the column of spherically-modeled gels at different rates, leading to their separation.
For downstream biopharmaceutical manufacturing, a thorough understanding of the phenomena of particle packing from a suspension is necessary to optimize efficiency and yield. The filtration of protein suspensions is an important step in isolating target proteins when purifying proteins in blood plasma for therapeutic use. These proteins may also be further processed using chromatography to further isolate the target protein with high purity and homogeneity. Modeling the interactions of the suspension mixtures and their packing behavior across a variety of pressure and flow systems provides an ideal methodology to optimize the separation of suspension mixtures. This, in turn, ensures efficiency in the transformation of valuable blood plasma to life-saving therapies. The modeling of the interactions in protein suspensions as well as their separations is computationally cost prohibitive today. Quantum computing provides an opportunity to provide insight and evaluate the separation process on a molecular level. The ability to both optimize the particle packing and model the interactions inherent in the separation could enable improved product yield and purity. However, given the limited capacities of today’s quantum computers, we focus on the second goal of optimizing the packing of the gels rather than modeling their interactions with the proteins.
When optimizing packing, the scale of the process is an additional complexity to consider. While it may be less computationally intensive to model smaller scale processes, it is commercially more valuable to understand how the interactions and packing dynamics change as the scale is increased. Economies of scale become increasingly important and ensuring that the efficiency can be achieved at larger and smaller scales is an important aspect of a wholistic understanding of the process. Therefore, while we only solve the problem today for a small, simple model of the problem, we provide a path to scaling up the algorithms applications to larger, more realistic models.
The utility of the packing is crucial with the ability to both maximize flux through the packed bed and increase opportunities for interaction with the bed being of great importance. During filtration, a desire to maximize the flux though the packing is required while for chromatography applications the primary goal is to maximize interactions with the packed bed. Both of these goals would be achieved by maximizing the tightness of the packing, providing great utility in the bio-manufacturing space. The ability to provide a size distribution profile of the various components of the suspension mixture is also key and would be an important determination in any implementation of the algorithm. It would enable the solution to be utilized across various implementations as well as provide a methodology to understand how particle size could affect the performance of the packed bed for either separations or chromatography. Thus, our solution allows for the exploration of an arbitrary number of distinct radii.
I.2 Sphere Packing
As mentioned above, we will focus on the packing aspect of protein chromatography. The gels can be modeled as spheres which are to be packed into a cylindrical column as tightly as possible, meaning in a way that minimizes empty space. This is the mathematical problem known as sphere packing. There are three features that can be toggled when discussing sphere packing: homogeneity, boundedness, and dimension: The spheres’ radii can be homogeneous (all the same) or heterogeneous (different). The problem can be unbounded (the packer has unlimited space available) or bounded (the packer has limited space available). Finally, the problem can be solved in two dimensions (circle packing) or three dimensions (sphere packing). While we consider both the homogeneous and heterogeneous case, we only consider the bounded case as chromatography is bounded by the filtration cylinder. Finally, while sphere packing can, in principle [1], be formulated in any dimension, we start with two dimensions for simplification and later give a path to the practical case of three dimensions.
Unbounded homogeneous packing has a long history. In 1773, Joseph Lagrange proved that the optimal circle packing is the hexagonal packing arrangement [2], which achieves a packing density (ratio of non-empty to empty space) of $\pi/(2\sqrt{3})\approx 0.91$. As for sphere packing, in 1611 Johannes Kepler conjectured that cubic close packing and the hexagonal close packing are optimal in their density of $\pi/(3\sqrt{2})\approx 0.74$. His conjecture was finally proved, with the aid of computer proof assistants in, 2017 [3]. Much less is known about unbounded heterogeneous packing except for special ratios of radii for two sphere sizes [4], leaving a general analytical solution an open problem within the field. The existence of general or analytical solutions, is summarized for unbounded packing in Table 1.
@@ I.2 Sphere Packing
Table 1: Unbounded homogeneous/heterogeneous sphere packing in two and three dimensions.  indicates the existence of a general analytical solution while ? indicates the existence of a partial analytical solution but unknown general solution.
@@
Table 2: Bounded homogeneous/heterogeneous sphere packing in two and three dimensions. ? indicates the existence of a partial analytical solution but unknown general solution.  indicates no analytical solutions, only numerical approximations.
@@ I.3 QAOA
The algorithm is based on the quantum adiabatic theorem which states that if a system starts in the ground state of an initial Hamiltonian and evolves slowly enough to a final Hamiltonian, the system will end in the ground state of said final Hamiltonian. This provides a useful computational tool for discrete optimization problems: simply start in the ground-state of an easy to solve Hamiltonian and slowly evolve to a final Hamiltonian which encodes the minimization problem. According to the quantum adiabatic theorem, the system will end in the ground state of this final Hamiltonian, which will encode the solution to the minimization problem (the state that minimizes the cost function). QAOA is a discretized version of the quantum adiabatic algorithm: its variational parameters are the coefficients of discrete time evolution terms which are tuned so the time evolution of the system approximates the continuous time evolution from the initial to the final Hamiltonian. The number of discrete time evolution terms is called the number of layers ($p$) of the QAOA circuit. It is known that, in the limit as $p$ approaches infinity, the state of the QAOA circuit approaches the optimal state [10]. The QAOA circuit takes the form
@@
where $H_{C}$ is the cost Hamiltonian (encoding the minimization problem), $H_{M}$ is the mixing Hamiltonian (encoding the easy-to-solve initial Hamiltonian), and |$\ket{\psi_{\text{init}}}$ is the ground state of the mixing Hamiltonian $H_{M}$.
@@ II Homogeneous Sphere Packing
Sphere packing is an inherently continuous problem in that the spheres may be placed anywhere in space. One way to approximate this continuity (as no computer has infinite memory) is to encode the position of each sphere $s=(x,y,z)$ into the bits or qubits of the hardware. This allows for a great deal of flexibility over the system. The downside is that the number of bits/qubits required to encode the properties scales exponentially with the desired precision. The formulation approaches the continuous case as the precision goes to infinity. The other way to approximate continuity is to discretize space into a grid of points, each of which is assigned a bit/qubit whose value (1 or 0) indicates the existence (or non-existence) of a sphere placed at that location. This is a less flexible formulation but one whose bit/qubit number scales polynomially with desired precision. Because of its amenability to QAOA, this second formulation is chosen.
@@ II.1 Formulation
To simplify the problem to a level that can be run on today’s hardware, instead of packing spheres into a cylinder, consider packing circles into a circle representing a two-dimensional cross-section of said cylinder. We can then formulate packing as maximizing the number of circles (each of radius $r$) that can be placed within the boundary circle of radius $R$ (the radius of the cylinder) such that none of the circles overlap? Note that since all the circles have the same radii, maximizing the number of circles packed is equivalent to maximizing packing density. This problem can be mapped to the problem of Maximum Independent Set (MIS) via the following construction: place an evenly spaced grid of nodes within the boundary circle and assign edges between all pairs of nodes such that circles centered at said nodes would overlap. The integer program for this formulations is
@@
where $\bar{x}_{v}=1-x_{v}$ and each indicator variable $x_{v}$ is 1 or 0 if there is or isn’t a circle centered on node $v$, respectively. The objective is the sum of inverses of indicator variables over all of the nodes, because this represents the number of nodes without circles centered on them, which should be minimized. The constraint is that the sum of the product of the indicator variables between every pair of nodes which share an edge must be zero. Note that this can only be satisfied if, for every pair of nodes which share an edge, both are zero. This implies that no two spheres may exist if they would overlap.
Figure 3 shows how the connectivity of the node graph increases with the discretization spacing $a$ for fixed radius $r$: as $a$ decreases, the nodes are placed closer together, allowing more fall within $2r$ of each other (the condition for an edge to exist). An example circle of radius $r$ is shown in the center of each graph for reference.
Figure 3:
Note that this two dimensional formulation may be extended to the full three dimensional formulation (packing spheres into a cylinder) using the same integer program (Eq. 2) by stacking the slices on top of one another and adding edges between pairs of points closer than $2r$. One might be concerned that the connectivity of the graphs for the three dimensional case grows faster than polynomial with the number of points. However, as shown in Section IV, this growth can be bounded by a low degree polynomial.
@@ II.2 Classical Solution
Figure 4:
Figure 5:
Since we have formulated our problem as MIS (Eq. 2), it can be approximately solved via classical computation. Figure 4 shows the maximum independent set found via classical computation for various spacings. It shows that as the discretization spacing $a$ decreases, the solver is able to pack larger numbers of circles $n$ within the boundary circle (corresponding to higher density circle packings). This is because, in the limit as $a$ goes to 0, the formulation approaches the continuous version of circle packing. Note that we have omitted the underlying graphs for the latter half of the examples for clarity of visualization.
Figure 5 shows how the number of packable circles scales with the discretization spacing $a$. It appears linear for this discretization spacing interval (though we know that it ultimately must asymptotically approach the maximum number of packable spheres). Therefore, we see that classically, because MIS is NP-hard, it takes an exponential increase in time to achieve a diminishing increase in the number of circles packed. However, for a quantum computer, the resources required to simulate this formulation (Subection IV) grow polynomially. Additionally, quantum computers have access to extra (potentially optimization-boosting) properties such as superposition, entanglement, and tunneling which take exponential time to simulate classically. Thus, we turn now to a quantum solution for homogeneous circle packing.
@@ II.3.1 Hamiltonian Formulation
To formulate MIS (Eq. 2) as a Hamiltonian, we make the following transformation from the indicator variable to Pauli spin matrices:
$x_{v}\to(I_{v}-Z_{v})/2$ (3)
@@
$H_{C}=\frac{1}{2}\sum_{v\in V}Z_{v}+\frac{\lambda}{4}\sum_{v,w\in E}(Z_{v}Z_{w}-Z_{v}-Z_{w})$ (4)
@@ II.3.2 Experiment Setup
The hardware selected for this experiment was IQM’s, 20-qubit, superconducting device, Garnet [11]. It was chosen, in large part, because of the grid connectivity of its qubit coupling map (Figure 6). It’s “missing corners” which prevent it from being a full 5 by 5 grid actually make it a natural fit for circle packing inside a larger boundary circle, as circles placed at said “missing corners” would extend beyond the boundary. Additionally, the “missing qubit” at the bottom right-hand corner that would have made the coupling map symmetric is also unnecessary for the problem as any packing of circles may be universally shifted any one direction until at least one of the circles touches the boundary. Therefore, without loss of generality, we start our packing by placing one circle at the location of the “missing qubit” along the boundary (Figure 7). Furthermore, the existence of this starting circle precludes the placement of circles at node directly above it and the node directly to its left. This can be seen when comparing the graph of Garnet’s qubit coupling map (Figure 6) and the graph for the experiment (Figure 7). This allows us to reduce the 21 qubits of the naive graph (Figure 6 plus the one “missing” node) to the 18 nodes shown in Figure 7. Here, circles have radius $r=1$ and the boundary circle has radius $R_{b}=4.2$. The graph corresponds to a discretization spacing of $a=\sqrt{2}$. Figure 8 shows that the optimal solution (the maximum independent set) for this discretization spacing is the placement of 12 spheres. This leads to a packing density (fraction of the boundary circle’s area filled by packing circles) of 0.680.68~{}0.680.68. It has been proven that the optimal packing density for 12 spheres 0.740.74~{}0.740.74 (which has a boundary circle of radius 4.024.02~{}4.024.02 [12].

### v2

@@ Abstract
Column chromatography is an important process in downstream biopharmaceutical manufacturing that enables high-selectivity separation of proteins through various modalities, such as affinity, ion exchange, hydrophobic interactions, or a combination of the aforementioned modes. Current mechanistic models of column chromatography typically abstract particle-level phenomena, in particular adsorption kinetics. A mechanistic model capable of incorporating particle-level phenomena would increase the value derived from mechanistic models. To this end, we model column chromatography via sphere packing, formulating three versions, each with increasing complexity. The first, homogeneous circle packing, is recast as maximum independent set and solved by the Quantum Approximate Optimization Algorithm on a quantum computer. The second, heterogeneous circle packing, is formulated as a graphical optimization problem and solved via classical simulations, accompanied by a road map to a quantum solution. An extension to the third, heterogeneous sphere packing, is formulated mathematically in a manner suitable to a quantum solution. Finally, detailed resource scaling is conducted to estimate the quantum resources required to simulate the most realistic model, providing a pathway to quantum advantage.
@@ I Introduction
As quantum computational power continues to grow rapidly [1], it is paramount that industries which rely heavily on computation start to plan today for tomorrow’s quantum advantage. One such industry is the biopharmaceutical industry, which uses massive amounts of classical computation to aid everything from drug discovery to operational optimization [2]. In this work, we investigate the utility of quantum computing to aid in the mechanistic modeling of column chromatography. Column chromatography is a primary separation method in downstream biopharmaceutical manufacturing[3, 4]. The modeling of chromatography columns has been utilized in accelerating process development and process scale-up as well as to maximize process productivity[4, 5, 6, 7]. Mechanistic models, which incorporate mathematical models of the transport and adsorption processes present in column chromatography, provide greater accuracy and precision in comparison to empirical models based mainly on statistical methods[4, 5, 6, 7]. Current approaches to mechanistic modeling typically abstract the particle-level adsorption kinetics through the use of generic isotherms[4, 6, 7].
Here, we begin the process of modeling particle-level phenomena in chromatography columns. It is optimal for the chromatography separation process to generate tightly packed chromatography columns without affecting the integrity of the individual particles. This problem is modeled as bounded sphere-packing, in which spheres are packed to minimize empty space. Bounded sphere packing is further cast as a graphical optimization problem amenable to quantum solutions.
@@
Figure 1: Three complexity levels of sphere packing for modeling chromatography: homogeneous circle packing, heterogeneous circle packing, and heterogeneous sphere packing, along with their respective difficulty levels for classical and quantum solutions over time.
We break the problem of sphere packing into three levels of increasing complexity: homogeneous circle packing, heterogeneous circle packing, and heterogeneous sphere packing (Figure 1). We solve example of homogeneous circle packing on real quantum hardware that exists today. Details of the hardware experiment include discussions on hyperparameter optimization, compilation, and noisy simulation. Evidence of parameter concentration is also included. Numerical results for heterogeneous circle packing via classical computation were executed, and the problem was formulated in a manner suitable for quantum computation. Finally, heterogeneous sphere packing is shown to be an extension of the heterogeneous circle packing formulation, and resource estimation is performed to estimate the quantum resources required to solve it. We will argue that while classical computers can approximately solve all three formulations today, their difficulty in doing so scales exponentially with increasing problem size and formulation complexity. And in contrast, while quantum computers can only solve the simplest formulation today, there is good evidence that in time they will be able to solve all three levels of complexity, each with only polynomially scaling difficulty.
@@
Figure 2: Schematic of chromatography for protein filtration. Different proteins filter through the column of spherically modeled gels at different rates, leading to their separation.
I.1 Sphere Packing
@@
Table 1: Unbounded homogeneous/heterogeneous sphere packing in two and three dimensions.  indicates the existence of a general analytical solution while ? indicates the existence of only partial analytical solutions.
@@
Table 2: Bounded homogeneous/heterogeneous sphere packing in two and three dimensions. ? indicates the existence of only partial analytical solutions.  indicates no analytical solutions.
As mentioned above, we will focus on the packing aspect of protein chromatography. The chromatography resin can be modeled as porous spheres which are to be packed into a cylindrical column as tightly as possible, meaning in a way that minimizes empty space (Figure 2). This is the mathematical problem known as sphere packing. There are three features that can be toggled when discussing sphere packing: homogeneity, boundedness, and dimension: The spheres’ radii can be homogeneous (all the same) or heterogeneous (different). The problem can be unbounded (the packer has unlimited space available) or bounded (the packer has limited space available). Finally, the problem can be solved in two dimensions (circle packing) or three dimensions (sphere packing). While we consider both homogeneous and heterogeneous cases, we only consider bounded cases as chromatography is bounded by the filtration cylinder. Finally, while sphere packing can, in principle [8], be formulated in any dimension, we start with two dimensions for simplification and later give a path to the practical case of three dimensions.
Unbounded homogeneous packing has a long history. In 1773, Joseph Lagrange proved that the optimal circle packing is the hexagonal packing arrangement [9], which achieves a packing density of $\pi/(2\sqrt{3})\approx 0.91$. As for sphere packing, in 1611 Johannes Kepler conjectured that cubic close packing and the hexagonal close packing are optimal in their density of $\pi/(3\sqrt{2})\approx 0.74$. His conjecture was finally proven, with the aid of computer proof assistants, in 2017 [10]. Much less is known about unbounded heterogeneous packing except for special ratios of radii for two sphere sizes [11], leaving a general analytical solution an open problem within the field. The existence of general or analytical solutions is summarized for unbounded packing in Table 1.
@@ I.2 QAOA
The algorithm is based on the quantum adiabatic theorem which states that if a system starts in the ground state of an initial Hamiltonian and evolves slowly enough to a final Hamiltonian, then the system will end in the ground state of said final Hamiltonian. This provides a useful computational tool for discrete optimization problems: simply start in the ground-state of an easy to solve Hamiltonian and slowly evolve to a final Hamiltonian which encodes the minimization problem. According to the quantum adiabatic theorem, the system will end in the ground state of this final Hamiltonian, which will encode the solution to the minimization problem (the state that minimizes the cost function). QAOA is a discretized version of the quantum adiabatic algorithm: its variational parameters are the coefficients of discrete time evolution terms which are tuned so the time evolution of the system approximates the continuous time evolution from the initial to the final Hamiltonian. The number of discrete time evolution terms is called the number of layers ($p$) of the QAOA circuit. It is known that, in the limit as $p$ approaches infinity, the final state of the QAOA circuit approaches the optimal state of the problem [17]. The QAOA circuit takes the form
@@
where $H_{C}$ is the cost Hamiltonian (encoding the minimization problem), $H_{M}$ is a mixing Hamiltonian (encoding an easy-to-solve initial Hamiltonian), and |$\ket{\psi_{\text{init}}}$ is the ground state of the mixing Hamiltonian $H_{M}$.
@@ II Homogeneous Sphere Packing
Sphere packing is an inherently continuous problem in that the spheres may be placed anywhere in space. One way to approximate this continuity (as no computer has infinite memory) is to encode the position of each sphere $s=(x,y,z)$ into the bits or qubits of the hardware. This allows for a great deal of flexibility over the system. The downside is that the number of bits/qubits required to encode the properties scales exponentially with the desired precision. The formulation approaches the continuous case as the precision goes to infinity. The other way to approximate continuity is to discretize space into a grid of points, each of which is assigned a bit/qubit whose value (1 or 0) indicates the existence (or non-existence) of a sphere placed at that location, respectively. This is a less flexible formulation but one whose bit/qubit number scales only polynomially with desired precision. Because of its amenability to QAOA, this second formulation is chosen.
@@ II.1 Formulation
To simplify the problem to a level that can be run on today’s hardware, instead of packing spheres into a cylinder, consider packing circles into a circle representing a two-dimensional cross-section of said cylinder. We can then formulate packing as maximizing the number of circles (each of radius $r$) that can be placed within the boundary circle of radius $R_{b}$ (the radius of the cylinder) such that none of the circles overlap. Note that since all the circles have the same radii, maximizing the number of circles packed is equivalent to maximizing packing density. This problem can be mapped to the problem of Maximum Independent Set (MIS) via the following construction: place an evenly spaced grid of nodes $V$ within the boundary circle and assign edges $v,w\in E$ between all pairs of nodes $v,w\in V^{2}$ such that circles centered at them would overlap. The integer program for this formulations is
@@
where $\bar{x}_{v}=1-x_{v}$ and each indicator variable $x_{v}$ is 1 or 0 if there is or isn’t a circle centered on node $v$, respectively. The objective is the sum of the inverses of the indicator variables over all of the nodes. This represents the number of nodes without circles centered on them, which should be minimized. The constraint encodes that the sum of the product of the indicator variables between every pair of nodes which share an edge must be zero. Note that this can only be satisfied if, for every pair of nodes which share an edge, both are zero. This implies that no two spheres may exist if they would overlap.
Figure 3: Graphs for discretization spacings $a=$ 1.4, 1.1, 0.9 and 0.75. A red circle of radius $r=1$ is shown in the center of each boundary circle for reference.
Figure 3 shows how, for fixed radius $r$, the connectivity of the graph increases with decreasing discretization spacing $a$, the spacing between the nodes. As $a$ decreases, the nodes are placed closer together, allowing more of them to fall within $2r$ of each other, at which point they would overlap, triggering the existence of an edge.
Note that this two dimensional formulation may be extended to the full three dimensional formulation (packing spheres into a cylinder) using the same integer program (Eq. 2) by stacking these two-dimensional slices on top of one another and adding additional edges between pairs of points from different slices that are closer than $2r$ to each other. One might be concerned that the connectivity of the graphs for the three dimensional case grows faster than polynomial with the number of points. However, as shown in Section IV, this growth can be bounded by a low degree polynomial.
@@ II.2 Classical Solution
Figure 4: Schematic showing how the size $n$ of the maximum independent set of the graphs increases with decreasing spacing $a$. Note that the underlying graphs have been omitted for the latter half of the examples for clarity of visualization.
Figure 5: Graph showing that the number of circles packed generally increases with decreasing spacing.
As we have formulated our problem as MIS (Eq. 2), it can now be approximately solved via classical computation. Figure 4 shows the maximum independent set found via classical computation for various spacings. It indicates that as the discretization spacing $a$ decreases, the solver is able to pack larger numbers of circles $n$ within the boundary circle (corresponding to higher packing densities). This is because, in the limit as $a$ goes to 0, the formulation approaches the continuous version of circle packing.
Figure 5 shows how the number of packed circles scales with the discretization spacing $a$. It appears linear for this discretization spacing interval, though we know that it ultimately must asymptotically approach the maximum number of packable spheres. Therefore, we see that classically, because MIS is NP-hard, it takes an exponential increase in time to achieve a diminishing increase in the number of circles packed. However, for a quantum computer, the resources required to simulate this formulation (Subection IV) only grow polynomially.
@@ II.3.1 Hamiltonian Formulation
To formulate MIS (Eq. 2) as a Hamiltonian, we make the following transformation from the indicator variable to Pauli spin matrices
$x_{v}\to(I_{v}-Z_{v})/2,$ (3)
@@
$H_{C}=\frac{1}{2}\sum_{v\in V}Z_{v}+\frac{\lambda}{4}\sum_{v,w\in E}(Z_{v}Z_{w}-Z_{v}-Z_{w}),$ (4)
@@ II.3.2 Experiment Setup
The hardware selected for this experiment was IQM’s 20-qubit superconducting device, Garnet [18]. It was chosen, in large part, because of the grid connectivity of its qubit coupling map (Figure 6). It’s “missing corners” which prevent it from being a full 5 by 5 grid actually make it a natural fit for circle packing inside a larger boundary circle, as circles placed at said “missing corners” would extend beyond the boundary. Additionally, the “missing qubit” at the bottom right-hand corner that would have made the coupling map symmetric is also unnecessary for the problem as any packing of circles may be universally shifted any one direction until at least one of the circles touches the boundary. Therefore, without loss of generality, we start our packing by placing one circle at the location of the “missing qubit” along the boundary (Figure 7). Furthermore, the existence of this starting circle precludes the placement of circles at the node directly above it and the node directly to its left. This can be seen when comparing the graph of Garnet’s qubit coupling map (Figure 6) and the graph for the experiment (Figure 7). This allows us to reduce the 21 qubits of the naive graph (Figure 6 plus the one “missing” node) to the 18 nodes shown in Figure 7. Here, circles have radius $r=1$ and the boundary circle has radius $R_{b}=4.2$. The graph corresponds to a discretization spacing of $a=\sqrt{2}$. Figure 8 shows that the optimal solution (the maximum independent set) for this discretization spacing is a placement of 12 spheres. This leads to a packing density (fraction of the boundary circle’s area filled by packing circles) of 0.680.68~{}0.680.68. It has been proven that the optimal packing density for 12 spheres 0.740.74~{}0.740.74 (which has a boundary circle of radius 4.024.02~{}4.024.02 [19]).